### v2

# Everything Everywhere All At Once: A First Principles Numerical Demonstration of Emergent Decoherent Histories

Abstract
Within the histories formalism the decoherence functional is a formal tool to investigate the emergence of classicality in isolated quantum systems, yet an explicit evaluation of it from first principles has not been reported. We provide such an evaluation for up to five-time histories based on exact numerical diagonalization of the Schrödinger equation. We find a robust emergence of decoherence for slow and coarse observables of a generic random matrix model and extract a finite size scaling law by varying the Hilbert space dimension over four orders of magnitude. Specifically, we conjecture and observe an exponential suppression of coherent effects as a function of the particle number of the system. This suggests a solution to the preferred basis problem of the many worlds interpretation (or the set selection problem of the histories formalism) within a minimal theoretical framework—without relying on environmentally induced decoherence, quantum Darwinism, Markov approximations or ensemble averages. We further discuss the implications of our results for the wave function of the Universe, interpretations of quantum mechanics and the arrow(s) of time.
I Introduction
Do we live in a quantum Multiverse? Since the formulation of the many worlds interpretation (MWI) Everett (1957); Witt (1970); Vaidman (2021), this idea enjoys increasing popularity among researchers Carr (2007); Saunders et al. (2010); Wallace (2012), popular science Carroll (2022) and pop culture EEA (2022). Although the basic premise of the MWI is “simply” to assume unitary evolution for the entire Universe (including its observers), its seemingly absurd consequence that the Universe consists of many universes existing in parallel (the Multiverse) is a source of strong controversies Gisin (2022). Using non-relativistic quantum mechanics, our contribution is to provide direct evidence—within a minimal first principles setup—that the MWI is compatible with our experienced “classical reality”.
It is useful to clarify from the start that the idea of the quantum Multiverse and the MWI should be distinguished from other potential Multiverses based on, e.g., an infinitely large Universe, many inflationary bubbles, different fundamental constants or various solutions in the string theory landscape Tegmark (2003); Carr (2007). Those universes are thought of being separated in spacetime, and thus admit in principle a classical understanding. In contrast, the Multiverse of the MWI consists of different universes at the same spacetime location.
One essential technical problem associated to the MWI is called the preferred basis problem: how to reconcile the Multiverse with our perceived classical experience within one universe? If Schrödinger’s equation is applied to the entire Universe, then by linearity superpositions proliferate and spread, splitting the wave function into many universes, or histories, evolving in parallel (also called branches, worlds, realities, narratives, etc.). However, the wave function can—without approximation—be split with respect to many different bases (indeed, a continuum of bases) and each basis provides a priori an equally justified starting point (and as we will discuss the wave function can be also split backwards in time). But as Bohr, in endless discussions with Einstein about the double slit experiment (and others), has made clear, a state describing a superposition of different properties makes said property ontologically indeterminate: in the double slit experiment the particle has no meaningful location without measurement Barad (2007). Thus, without identifying an additional structure justifying a classical description, as we experience it, the MWI describes infinitely many ontologically indeterminate splittings: a priori none of them allows to speak about “histories”, “worlds” or “realities” in any conventional, meaningful sense.
This additional structure must be derived within the MWI, and within non-relativistic quantum mechanics a satisfactory derivation must comply at least with two minimal desiderata. (A) The system is isolated, evolves unitarily and is prepared in a pure state. This avoids the introduction of any form of classical noise from the outside (e.g., in form of ensemble averages), which potentially implies the answer to the question already from the start. We remark that the system might be (but does not need be) split into subsystems. (B) The condition of classicality must be a meaningful, rigorous definition that is suitable to account for multi-time properties or temporal correlation functions because the perception of classicality is a repeated experience (adapting Einstein’s quote, the moon was there yesterday, is there today and will be there tomorrow). This is important: speaking of different worlds or histories becomes meaningful if we can reason about their past, present and future in classical terms.
Here, we use the decoherence functional (DF) introduced within the consistent or decoherent histories framework (or simply “histories framework” for short) Griffiths (1984); Gell-Mann and Hartle (1990); Omnès (1992); Dowker and Halliwell (1992); Gell-Mann and Hartle (1993); Halliwell (1995); Dowker and Kent (1996); Griffiths (2002); Gell-Mann and Hartle (2007); Griffiths (2019) to rigorously investigate the emergence of multi-time classicality (see Sec. II for the technical details). We numerically show that bases defined by a slow and coarse observable of a non-integrable system are robustly decoherent according to this criterion. More specifically, we show for up to five-time histories and a Hilbert space dimension $D$ varying over four orders of magnitude that quantum effects (suitably quantified below) are exponentially suppressed as a function of the particle number $N$ of the system (with $N\sim\log D$). This provides a firm starting point to discuss the MWI and to address the set selection problem of the histories framework Paz and Zurek (1993); Dowker and Kent (1995, 1996); Riedel et al. (2016). We further discuss important cosmological implications, namely that the emergence of classicality is ubiquituous, that almost all initial wave functions can give rise to interesting universes, and that the branching of the wave function is a priori not related to any arrow of time.
Perhaps surprisingly, our results do not rely on environmentally induced decoherence (EID) Zurek (2003); Joos et al. (2003); Schlosshauer (2019) [^1] Note the unfortunate double meaning of decoherence in the literature:a priori there is no connection between the DF in the histories framework and the concept of EID. or its refinement to quantum Darwinism Zurek (2009); Korbicz (2021); Zurek (2022). This is surprising because the widely proclaimed (sole) answer to the question why the MWI gives rise to classically looking universes is EID, and many researchers worked on connecting the MWI or the histories framework to EID and quantum Darwinism Finkelstein (1993); Saunders (1993); Paz and Zurek (1993); Diósi et al. (1995); Brun (1997); Yu (1998); Vaidman (1998); Brun (2000); Riedel et al. (2016); Albrecht et al. (2022); Touil et al. (2022); Strasberg (2023), yet an explicit evaluation of the DF following the desiderata (A) and (B) is still missing. Only if one invokes a rigorous notion of multi-time quantum Markovianity Pollock et al. (2018); Li et al. (2018); Milz and Modi (2021), clear connections between the DF and EID have been established Paz and Zurek (1993); Diósi et al. (1995); Brun (1997); Yu (1998); Brun (2000); Strasberg (2023), but this only shifts the problem of proving multi-time decoherence to proving multi-time Markovianity, which is a daunting task too Dümcke (1983); Ford and O’Connell (1996); Figueroa-Romero et al. (2019, 2021); Strasberg et al. (2023).
In contrast, the present approach does not rely on any system-environment tensor product splitting, although it is important to emphasize that it is not in conflict with EID or quantum Darwinism when applied to it. Instead, we only consider slow and coarse observables of isolated, non-integrable quantum systems, similar to the approach taken by van Kampen in 1954 Van Kampen (1954). While the importance of slow and coarse (or quasi-conserved) observables for the emergence of decoherence has been anticipated in the histories formalism Gell-Mann and Hartle (1993); Brun and Halliwell (1996); Halliwell (1998, 1999a); Calzetta and Hu (1999); Halliwell (2003); Gell-Mann and Hartle (2007); Halliwell (2010), non-integrability has never been considered a key factor by proponents of the histories, EID or quantum Darwinism framework: van Kampen’s work remained ignored despite him emphasizing its importance for the quantum-to-classical transition van Kampen et al. (2000); van Kampen (2008). Only recently rigorous evidence for van Kampen’s ideas slowly accumulated Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Strasberg et al. (2023); Strasberg (2023). Still, it has not yet been considered in light of the MWI and a first principles evaluation of the DF remains missing.
To summarize, our objective is to subject the MWI and the question whether it is compatible with our perceived classical reality to a rigorous, quantitative test based on the minimal desiderata (A) and (B). To this end, we introduce the general theoretical framework to address this question in Sec. II. Then, we present extensive numerical results for a heat exchange model (an archetypical example of a nonequilibrium process in thermodynamics) in Sec. III. Section IV discusses important implications and we conclude and provide perpectives in Sec. V.
II General Framework
II.1 Mathematical definitions and problem
We consider an isolated quantum system with Hilbert space ${\cal{H}}$ of dimension $D=\dim{\cal{H}}$ and Hamiltonian $H$. The time evolution operator from $t_{j}$ to $t_{k}$ is denoted $U_{k,j}=e^{-iH(t_{k}-t_{j})}$ ($\hbar\equiv 1$) and the initial state is $|\psi(t_{0})\rangle$. Furthermore, $\{\Pi_{x}\}_{x=1}^{M}$ denotes a complete set of $M$ orthogonal projectors satisfying $\sum_{x=1}^{M}\Pi_{x}=I$ (with $I$ the identity) and $\Pi_{x}\Pi_{y}=\delta_{x,y}\Pi_{x}$. They divide the Hilbert space into a direct sum of subspaces: ${\cal{H}}=\bigoplus_{x=1}^{M}{\cal{H}}_{x}$. The dimension of the subspace ${\cal{H}}_{x}$ equals the rank of the projector $\Pi_{x}$ and is denoted $V_{x}=\mbox{tr}\{\Pi_{x}\}=\dim{\cal{H}}_{x}$. Note that $D=\sum_{x=1}^{M}V_{x}$.
Next, we decompose the unitarily evolved state $|\psi(t_{n})\rangle=U_{n,0}|\psi(t_{0})\rangle$ by writing $U_{n,0}=U_{n,n-1}\cdots U_{1,0}$ and inserting identities at times $t_{n}>\dots>t_{1}>t_{0}$:
$|\psi(t_{n})\rangle$ $=\sum_{x_{n}}\Pi_{x_{n}}U_{n,n-1}\cdots\sum_{x_{1}}\Pi_{x_{1}}U_{1,0}\sum_{x_{0}}\Pi_{x_{0}}|\psi(t_{0})\rangle$
$|\psi(\textbf{x})\rangle.$ x ) ⟩ . (1)
Here, we abbreviated $\textbf{x}=(x_{n},\dots,x_{1},x_{0})$, which we call a history of length $L=n+1$ in the following. Moreover, $|\psi(\textbf{x})\rangle=\Pi_{x_{n}}U_{n,n-1}\cdots\Pi_{x_{1}}U_{1,0}\Pi_{x_{0}}|\psi(t_{0})\rangle$ is the (non-normalized) state conditional on “passing through” subspaces ${\cal{H}}_{x_{j}}$ at times $t_{j}$. Note that this construction becomes identical to Feynman’s path integral if we let $t_{j+1}-t_{j}\rightarrow 0$ and consider projectors $\Pi_{x}$ in the position representation.
Finally, we introduce the decoherence functional
${\mathfrak{D}}(\textbf{x};\textbf{y})\equiv{\langle{\psi(\textbf{y})|\psi(\textbf{x})}\rangle},$ (2)
which is a Hermitian $M^{L}\times M^{L}$ matrix. Moreover, the important decoherent histories condition (DHC) is defined by the condition
[…]
The central task of this work is to understand the conditions when the DHC is generically satisfied. In particular, we will see that it is hard to strictly satisfy Eq. (3) for the situations we are interested in. Hence, we will quantitatively study from first principles a suitable smallness condition ${\mathfrak{D}}(\textbf{x};\textbf{y})\approx 0$ discussed in Sec. II.3 below. We remark that this is a well-defined and non-trivial mathematical problem—independent of the physical meaning attached to the different objects to which we turn now.
II.2 Physical meaning and further terminology
We continue with physical clarifications, also related to the previous literature.
First, we call $\{\Pi_{x}\}_{x=1}^{M}$ a coarse-graining and $x$ a macrostate, which is conventional terminology in statistical mechanics. Indeed, we are interested in observables that humans can perceive, and those are necessarily coarse. Quantitatively, this means that the number of projectors is much smaller than the Hilbert space dimension: $M\ll D$. Note that this is also satisfied in any quantum experiment that relies on macroscopic detectors.
Next, we briefly clarify the meaning of the DHC, which has been already discussed in the literature Griffiths (1984); Gell-Mann and Hartle (1990); Omnès (1992); Dowker and Halliwell (1992); Gell-Mann and Hartle (1993); Halliwell (1995); Dowker and Kent (1996); Griffiths (2002); Gell-Mann and Hartle (2007); Griffiths (2019). [^2] There has also been some controversy about the precise mathematical formulation of the DHC (see, e.g., Ref. Diósi (2004)), but it seems that Eq. (3) is nowadays universally accepted Griffiths (2019). To this end, recall that according to our classical (Newtonian or pre-quantum) ontology the world out there is made up of entities with independent and well-defined properties that can be revealed in principle to arbitary precision, i.e., any uncertainty about the state of the world is entirely subjective or epistemic (for a broader discussion see, e.g., Ref. Barad (2007)). This worldview is challenged by quantum physics. However, the DHC guarantees that, for the process describing the macrostates $x_{j}$ at times $t_{j}$, such a classical ontology becomes quantitatively accurate because Eq. (3) describes the absence of detectable quantum interference effects. This makes the dynamics of the coarse properties $x_{j}$ isomorphic to a classical stochastic process (see also Refs. Smirne et al. (2018); Strasberg and Díaz (2019); Milz et al. (2020a, b); Milz and Modi (2021); Strasberg (2022); Strasberg et al. (2023); Strasberg (2023); Szańkowski and Cywiński (2023)) and implies the validity of Leggett-Garg inequalities Emary et al. (2014).
Before turning to the connection with the MWI, it is important to be clear about the fact that the DHC has differing ontological statuses even among practitioners of the histories formalism Omnès (1992); Dowker and Kent (1996). For instance, Griffiths regards the DHC as primary, that is, quantum evolution is fundamentally stochastic and the deterministic Schrödinger equation can only be used to compute the DF if the DHC is satisfied Griffiths (1984, 2002, 2019). In contrast, we here take unitary evolution as primary (as in the MWI) and view the DHC as an emergent property. This means we use the histories formalism as a convenient mathematical tool to address a well-defined problem.
We deliberately point out that this attitude shall neither imply that the MWI is correct not that the consistent histories interpretation of Griffiths is incorrect. In contrast, the question we ask here sheds light on both interpretations. It is important to know for the MWI whether the quantum Multiverse supports decoherent histories, and likewise the question with which histories Schrödinger’s equation is compatible, and whether this compatibility is exact or approximate, influences the scope of the consistent histories interpretation.
We continue by commenting on the connection between the DHC and the MWI as we view it in this work. As emphasized in the introduction, the definition of a “world” within the MWI is a priori complicated by the existence of a continuum of mathematically conceivable different worlds (preferred basis problem). Moreover, even among the proponents of the MWI there is no universally agreed on quantitative definition of a world. Here, we use this liberty and interpret the DHC as a minimal criterion to define a world within the MWI. This is motivated by what we said above: the classical world that we perceive is compatible with quantum mechanics precisely when the observables we look at satisfy the DHC. We believe this view is in unison with proponents of the MWI, for instance, Vaidman verbally defines a world as “the totality of macroscopic objects […] in a definite classically described state” Vaidman (2021). Moreover, a close connection between the DHC and the existence of records has been established by various researchers Albrecht (1992); Gell-Mann and Hartle (1993); Finkelstein (1993); Paz and Zurek (1993); Halliwell (1999b); Dodd and Halliwell (2003); Riedel et al. (2016); Hartle (2016).
In the following, it is also advisable to drop the word “classical” in favor of “decoherent”, at least in a technical context. Indeed, depending on the context, classicality can have many different meanings since the boundary between quantum and classical physics can not be reduced to a single condition. Thus, in short, our terminology is the following: histories that satisfy exactly or approximately (see Sec. II.3) the DHC are called decoherent (and only sometimes classical) and in the context of the MWI we call those histories also branches or worlds. We also repeat once more that the DHC is not identical to but compatible with the concept of EID.
We conclude by specifying what we mean by a first principles demonstration of the DHC. Essentially, we aim at a general understanding of the DHC by using physical assumptions that are not in conflict with the framework of Sec. II.1. This has two major implications. First, we do not allow for any assumption that breaks unitarity and, instead, we solve the Schrödinger equation exactly. In particular, we do not use Markov approximations as done in Refs. Paz and Zurek (1993); Diósi et al. (1995); Brun (1997); Yu (1998); Brun (2000); Smirne et al. (2018); Strasberg and Díaz (2019); Milz et al. (2020b); Strasberg (2023), which—even though insightful—shift the problem from justifying classicality to justifying Markovianity.
Second, we demand that the state of the isolated system is pure. By avoiding classical ensemble averages, any remaining uncertainty must stem from the quantum state or dynamics. Indeed, the only explicit evaluation of the DHC for a non-trivial many-particle system (and not relying on Markov approximations) has been done using the influence functional for a harmonic oscillator environment (Caldeira-Leggett model), assuming that the environment is prepared in a canonical Gibbs ensemble Schmid (1987); Dowker and Halliwell (1992); Gell-Mann and Hartle (1993); Halliwell (1999b, 2001); Subaşı and Hu (2012). This is problematic as it remains unclear whether the observed decoherence is a consequence of the dynamics itself (as one would hope for) or a consequence of the classical ensemble average.
What remains beyond Markov approximations and ensemble averages, and beyond conserved quantities that trivially satisfy the DHC, are indirect arguments based, e.g., on the approximate behavior of projectors in the Heisenberg picture Omnès (1989, 1992); Gell-Mann and Hartle (1993); Brun and Halliwell (1996); Halliwell (1998, 1999a); Calzetta and Hu (1999); Halliwell (2003); Gell-Mann and Hartle (2007); Halliwell (2010). [^3] It is also worth to note that most of these indirect arguments aimed at establishing the DHC plus the fact that there is only a single history happening (i.e., a deterministic history framework). This is more than what is necessary in our view (the existence of classical stochastic processes is an experimental fact). Moreover, somewhat confusingly this stronger condition (DHC plus determinism) has been often called “quasi-classicality”, which verbally suggests something weaker than classicality as used here. This is another reason for us to prefer the word decoherence over the word classicality. But an explicit evaluation of the DHC or even an estimate of it has never been attempted therein. It is this important gap that we fill with the present work, which is backed up by the general idea of van Kampen Van Kampen (1954) (see also Sec. II.4) and supported only by a few preliminary results so far Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Strasberg et al. (2023); Strasberg (2023).
II.3 Approximate decoherence
For reasons that will become clear below, exact decoherence, i.e., a strict satisfaction of Eq. (3), is not the rule. All we can realistically hope for is to satisfy the DHC approximately, even though our results will also indicate that this approximation is typically so good that it becomes indistinguishable from exact decoherence for all practical purposes.
Approximate decoherence is typically quantified—after realizing $|{\mathfrak{D}}(\textbf{x};\textbf{y})|\leq\sqrt{{\mathfrak{D}}(\textbf{x};\textbf{x}){\mathfrak{D}}(\textbf{y};\textbf{y})}$ due to Cauchy-Schwarz—by introducing a “normalized DF” and demanding that Dowker and Halliwell (1992)
$\epsilon(\textbf{x};\textbf{y})\equiv\frac{|{\mathfrak{D}}(\textbf{x};\textbf{y})|}{\sqrt{{\mathfrak{D}}(\textbf{x};\textbf{x}){\mathfrak{D}}(\textbf{y};\textbf{y})}}\ll 1\text{ for all }\textbf{x}\neq\textbf{y}.$ (4)
[…]
We therefore choose a dual strategy to quantify decoherence in this work. First, we consider the average of the non-trivial values of $\epsilon(\textbf{x};\textbf{y})$, defined as
$\epsilon\equiv\frac{1}{M^{2L-1}-M^{L}}\sum_{\textbf{x}\neq\textbf{y}}\epsilon(\textbf{x};\textbf{y}).$ (5)
This is probably the simplest quantifier one can consider, but we believe its simplicity makes it appealing to get a first impression of what is going on.
Figure 1: Example illustrating the DF for a coarse-graining with six macrostates and eight time steps. Two possibly interfering histories are indicated by dashed and dotted lines (the lines are used for visualization only, within this DF only their macrostates at times $t_{i}$ are defined).
However, to rigorously quantify the worst case scenario in an operationally meaningful way, we also consider a second strategy. To this end, we select an arbitrary subset $T\subset\{t_{0},t_{1},\dots,t_{n}\}$ of times on which the DF is defined, for instance, $T=\{t_{2},t_{7}\}$ in the sketch of Fig. 1. Furthermore, we denote by $Z_{T}$ all possible histories z that we can associate to $T$ with respect to the given coarse-graining, e.g., with respect to Fig. 1 these are all 2-time histories $Z_{T}=\{x^{1}_{7},x^{2}_{7},\dots,x^{6}_{7}\}\times\{x^{1}_{2},x^{2}_{2},\dots,x^{6}_{2}\}$ (with $\times$ the Cartesian product).
Next, we construct two types of probabilities on $Z_{T}$. The first is the actual Born rule probability to measure $z\in Z_{T}$, which we can get from the DF defined on $T$ via
$p(\textbf{z})\equiv\sum_{\textbf{x}(\textbf{z}),\textbf{y}(\textbf{z})}{\mathfrak{D}}(\textbf{x};\textbf{y}),$ (6)
where $\sum_{\textbf{x}(\textbf{z})}$ indicates a sum running over all x holding z fixed. For instance, for the example in Fig. 1 (the pink bars) we have $\sum_{\textbf{x}(\textbf{z})}=\sum_{\textbf{x}}\delta_{x_{7},x^{3}_{7}}\delta_{x_{2},x_{2}^{4}}$ and thus
$p(\textbf{z})={\langle{\psi(t_{0})|U^{\dagger}_{2,0}\Pi_{x_{4}}U^{\dagger}_{7,2}\Pi_{x_{3}}U_{7,2}\Pi_{x_{4}}U_{2,0}|\psi(t_{0})}\rangle}.$ (7)
The second probability we can introduce is a decohered or classical version of $p(\textbf{z})$, obtained under the assumption that the DHC is obeyed. It is defined as a sum over the diagonal elements only:
$p_{\text{cl}}(\textbf{z})\equiv\sum_{\textbf{x}(\textbf{z})}{\mathfrak{D}}(\textbf{x};\textbf{x}).$ (8)
Now, to evaluate the difference between the true quantum probability distribution $p(\textbf{z})$ and its decohered counterpart $p_{\text{cl}}(\textbf{z})$ we use the $L_{1}$-norm or trace distance
$\Delta_{T}(p|p_{\text{cl}})\equiv\frac{1}{2}\sum_{\textbf{z}\in Z_{T}}|p(\textbf{z})-p_{\text{cl}}(\textbf{z})|\in[0,1].$ (9)
Inserting the definitions, we see that the trace distance is determined by the off-diagonal elements of the DF:
$\Delta_{T}(p|p_{\text{cl}})=\frac{1}{2}\sum_{\textbf{z}\in Z_{T}}\left|\sum_{\textbf{x}(\textbf{z})\neq\textbf{y}(\textbf{z})}{\mathfrak{D}}(\textbf{x};\textbf{y})\right|.$ (10)
In addition, the trace distance has a clear operational meaning in a hypothesis testing scenario Wilde (2019). Namely, it determines in an optimized single shot scenario the minimum probability $P_{\text{min}}(\text{fail})=[1-\Delta_{T}(p|p_{\text{cl}})]/2$ for an experimenter to fail to distinguish between the coherent and decohered quantum process, described by $p(\textbf{z})$ and $p_{\text{cl}}(\textbf{z})$, respectively. For a given measurement resolution this gives the desired criterion to decide when a set of approximately decoherent histories is “decohered enough”.
However, the trace distance (9) still depends on $T$, but we are only interested in the worst case scenario. Moreover, we notice the following fact: if $t_{n}\notin T$, then the sum over $x_{n}$ and $y_{n}$ does not contribute to Eq. (9). This follows from a containment property of the DF: the DF for histories of length $L^{\prime}<L$ can be obtained from the DF for histories of length $L$ by tracing out the final $L-L^{\prime}$ time steps. Thus, we will plot in Sec. III the maximum
$\Delta^{\text{max}}_{L}\equiv\max_{T\cap\{t_{n}\}\neq\emptyset}\Delta_{T}(p|p_{\text{cl}}).$ (11)
Note that this number still depends on $L$.
II.4 Physical origin of decoherence
The physical origin of decoherence has been discussed at many places and shall not be repeated here in detail. We therefore limit ourselves to listing the four assumptions used by van Kampen Van Kampen (1954) (see also Refs. Strasberg et al. (2023); Strasberg (2023) for a more recent discussion) followed by a perhaps oversimplified yet hopefully intuitive analogy.
The four assumptions are the following. First, one needs a large Hilbert space dimension $D$, which is certainly satisfied for macroscopic systems. Typically, $D={\cal{O}}(10^{N})$ and $N={\cal{O}}(10^{23})$ to give some numbers. Second, one must consider a coarse coarse-graining satisfying $M\ll D$. As discussed above, this is satisfied for all that we humans can perceive. Third, the considered projectors must evolve slowly in a suitable sense such that the unobserved microscopic degrees of freedom have time to self-average or randomize. Common examples include hydrodynamic modes, collective degrees of freedom, weakly coupled subsystems, or more generally projectors that almost commute with the total Hamiltonian. Fourth, the isolated system should be non-integrable.
This last point is certainly still subject to debate, not the least because a universally agreed on definition of quantum (non-)integrability is lacking. For instance, van Kampen assumed that the energy gap spectrum of the Hamiltonian $H$ is non-degenerate (apart from rare accidental degeneracies) Van Kampen (1954), the results of Ref. Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Strasberg et al. (2023) were based on the eigenstate thermalization hypothesis Srednicki (1999); D’Alessio et al. (2016); Deutsch (2018), and here we use random matrix theory as also done in Ref. Nation and Porras (2020); Strasberg (2023). Since already classical systems tend to be chaotic (e.g., the Newtonian three-body problem), we believe that non-integrability can hardly count as an assumption in our Universe, even though it is an open question whether it is strictly necessary.
Intuitively, we like to motivate the need for non-integrability by the following analogy. Since the DHC describes the lack of interference between different macrostates $x$, we like to picture the coherences of all microscopic degrees of freedom as ripples caused by stones thrown into an initially still pond. If one zooms in very much (corresponding to very large $M$) or considers only a single stone thrown into the pond (corresponding to very small $D$), one can certainly see a clear wave pattern and interference effects. But if one zooms out a bit, or averages over a small area of the pond, and throws in a lot of stones (at different places), it becomes very hard to see any significant interference effects. One reason for it is simply the law of large numbers: there are many more combinations possible where ripples from different stones destructively interfere than possible combinations of constructive interference. This is identical to rolling $10^{23}$ dice: the vast majority of sequences has an average face value around 3.5 and the number of sequences deviating significanty from it is exponentially suppressed.
However, there is also another effect at work. In reality all stones have a slightly different size and weight, thus causing ripples with, e.g., different amplitudes or velocities. This implies that even if at a certain moment in time the state of the stones is highly synchronized (for instance, they could have been thrown symmetrically into the pond), this synchronized state will very quickly dephase and look generic. This effect is absent in integrable systems where the extensive amount of conserved quantities causes many regularities in the dynamics. It is thus reasonable to conjecture that non-integrabe systems, here corresponding to stones of different sizes and weights, show a more robust emergence of decoherence for a broader class of initial states.
III Numerical Demonstration
III.1 Setup
We will now explicitly validate the idea that coarse and slow observables of a non-integrable many-body system satisfy the DHC with increasing accuracy for increasing Hilbert space dimension $D$. This is done by exact numerical diagonalization for a setup consisting of two identical, coupled systems exchanging energy. We choose this archetypical setup of a nonequilibrium thermodynamic process because it is intuitive and allows us to clearly identify arrow(s) of time associated with the flow of heat from hot to cold. In principle, this flow of heat could then be harnessed to create work or free energy, an important prerequisite for the formation of life and intelligent observers in the Universe, but (obviously) our code can not simulate any forms of life or observers.
Figure 2: Sketch of two identical, interacting systems $A$ and $B$ with discrete energy levels. The energies of both $A$ and $B$ are coarse-grained into windows $x$ with an increasing number of levels. Finally, the dynamics is restricted to a microcanonical subspace of the total energy corresponding to windows $(x_{A},x_{B})=(+1,-1)\cup(0,0)\cup(-1,+1)$.
We call our systems $A$ and $B$ (see Fig. 2 for a sketch) and the total Hamiltonian is
$H=H_{A}+H_{B}+\lambda H_{I},$ (12)
where $H_{I}$ is the interaction and $\lambda$ a small parameter to ensure weak coupling (see below). We coarse-grain the local energies of $A$ and $B$ by projectors $\Pi_{x,x^{\prime}}\equiv\Pi_{x}^{A}\otimes\Pi_{x^{\prime}}^{B}$ for a pair of integers $(x,x^{\prime})$. In units of some energy precision or width $\Delta E$, $\Pi_{x,x^{\prime}}$ projects on the energy interval $[x,x+1)\times[x^{\prime},x^{\prime}+1)$, i.e., $\Pi_{x}^{A}$ is spanned by all energy eigenstates $|k\rangle_{A}$ of $H_{A}$ whose energy eigenvalue $E_{k}$ satisfies $x\leq E_{k}/\Delta E<x+1$, and similarly for $\Pi_{x^{\prime}}^{B}$.
In the following, we consider one microcanonical subspace with some fixed total energy $E_{\text{tot}}=E_{A}+E_{B}$, where $E_{A}$ ($E_{B}$) is the coarse-grained energy of system $A$ ($B$) determined with precision $\Delta E$. Thus, suppose we fix the total energy to be $E_{\text{tot}}=m\Delta E$ with $m\in\mathbb{Z}$, then the projector on this microcanonical subspace is given by
$P_{m}=\sum_{x,x^{\prime}}\delta_{m,x+x^{\prime}}\Pi_{x,x^{\prime}}.$ (13)
The Hamiltonian restricted to this subspace is consequently given by
$H_{m}\equiv P_{m}HP_{m}=\sum_{x,y}\Pi_{x,m-x}H\Pi_{y,m-y}.$ (14)
Specifically, we set (without loss of generality) in the following $E_{\text{tot}}=0$. Moreover, as sketched in Fig. 2, we consider for simplicity only three participating energy macrostates in $A$ and $B$, i.e., $P_{0}=\Pi_{-1,+1}+\Pi_{0,0}+\Pi_{+1,-1}$. Of course, for a realistic macroscopic system more subspaces should be considered, but three subspaces are sufficient for a proof-of-principle demonstration of our main ideas.
Since the total energy is fixed, we write in the following $(\Pi_{-},\Pi_{0},\Pi_{+})\equiv(\Pi_{-1,+1},\Pi_{0,0},\Pi_{+1,-1})$ with associated subspace dimensions $(V_{-},V_{0},V_{+})$. Note that these projectors project on subspaces of systems $A$ and $B$. The restricted Hamiltonian (14) has consequently nine blocks:
$H_{0}=\begin{pmatrix}H_{--}&H_{-0}&H_{-+}\\ H_{0-}&H_{00}&H_{0+}\\ H_{+-}&H_{+0}&H_{++}\\ \end{pmatrix}.$ (15)
Each of these blocks is in principle fully determined by the microscopic Hamiltonian (12), but we only want to capture four main features here. First, the systems $A$ and $B$ are assumed to be identical from a thermodynamic point of view. In particular, we assume that the relation between energy and temperature is the same in $A$ and $B$ such that their energy difference is proportional to their temperature difference. Second, we consider normal thermodynamic systems where the subspace corresponding to an equal energy distribution (or the same temperature according to the previous agreement) is the largest “equilibrium” subspace, i.e., $V_{0}>V_{-},V_{+}$. Third, we assume a weak interaction between $A$ and $B$, i.e., the interaction energy $\lambda H_{I}$ is supposed to be negligible (otherwise it would not be justified to restrict the discussion to a microcanonical energy window defined by the sum $E_{A}+E_{B}$ of local energies only). Fourth, we want to mimic the interaction of two generic complex (and thus non-integrable) many-body systems. According to common knowledge in statistical mechanics, this can be efficiently done by using random matrix theory Wigner (1967); Brody et al. (1981); Deutsch (1991); Beenakker (1997); Guhr et al. (1998); Haake (2010); D’Alessio et al. (2016); Deutsch (2018); Reimann and Dabelow (2021).
In unison with these agreements we consider below the following specific Hamiltonian. The diagonal blocks $(H_{--},H_{00},H_{++})$ are modeled by diagonal matrices with ($V_{-},V_{0},V_{+}$) many equally spaced energies in the interval $[0,2\Delta E)$. [^4] The factor “2” comes from the fact that the energy uncertainty of the projector $\Pi_{x,x^{\prime}}\equiv\Pi_{x}^{A}\otimes\Pi_{x^{\prime}}^{B}$ is $2\Delta E$ because both $\Pi_{x}^{A}$ and $\Pi_{x^{\prime}}^{B}$ have an energy uncertainty $\Delta E$. Moreover, assuming $H_{--}$, $H_{00}$ and $H_{++}$ to be diagonal can always be achieved by a block-unitary transformation. Finally, equal energy spacing is used for convenience only, a random spacing, for instance, has been observed to give rise to the same behaviour. Moreover, we set $H_{--}=H_{++}$, implying $V_{-}=V_{+}$ such that the total Hilbert space dimension is $D=2V_{-}+V_{0}$. The coupling between different blocks is realized by random matrices $H_{-0}=H_{0-}^{\dagger}$ and $H_{0+}=H_{+0}^{\dagger}$ with elements drawn from an othogonal zero-mean-unit-variance Gaussian ensemble (the unitary Gaussian ensemble was not observed to give rise to different behaviour) multiplied by the small coupling constant $\lambda$. Note that $H_{-0}$ and $H_{0+}$ are not square matrices since $V_{0}>V_{-},V_{+}$. Moreover, in view of the weak interaction, we set $H_{-+}=H_{+-}^{\dagger}$ to be a matrix of zeros, i.e., we forbid transitions between energy levels that are too far away from each other.
To remain in the regime of weak coupling, but to ensure that the different energy levels in $A$ and $B$ also sufficiently interact to make energy transport efficient, we have to choose $\lambda$, $\Delta E$ and the volumes appropriately. Their values can be estimated as follows. First, let us define weak coupling by demanding that the energy of the diagonal part $H_{0}=H_{--}+H_{00}+H_{++}$ dominates the interaction energy $\lambda H_{I}=H_{-0}+H_{0-}+H_{+0}+H_{0+}$. To estimate their typical value we use an average over the microcanonical ensemble ${\langle{\dots}\rangle}_{\text{mic}}$. For the diagonal part we find the value ${\langle{H_{0}}\rangle}_{\text{mic}}\approx\Delta E$. Since ${\langle{\lambda H_{I}}\rangle}_{\text{mic}}=0$, we look at the standard deviation $\lambda{\langle{H_{I}^{2}}\rangle}_{\text{mic}}^{1/2}\approx\lambda\sqrt{V_{0}V_{\pm}/D}\approx\lambda\sqrt{V_{\pm}}$. Thus, we obtain the condition $\lambda^{2}V_{\pm}/\Delta E^{2}\ll 1$ (note that we take $\lambda$ to have the dimension of energy). To make $A$ and $B$ sufficiently interact we note that the interaction smears out the local energy eigenstates of $A$ and $B$ proportional to $\lambda$ (“level broadening”). At the same time, the density of states in the $x=\pm 1$ subspace is roughly $\Delta E/V_{\pm}$. To guarantee that the smeared out levels in $A$ ($B$) overlap with sufficiently many levels in $B$ ($A$), we thus demand $(\lambda V_{\pm}/\Delta E)^{2}\gg 1$. Specifically, what we found to work numerically well are the conditions
[…]
compare also with Ref. Bartsch et al. (2008) for an analytical study of a similar model (two equal energy bands coupled via a random matrix). Unless otherwise mentioned, we choose $\lambda$ below such that the left equation reduces to $0.01$ $(\ll 1)$, which implies that the right equation is well satisfied for $V_{\pm}\gtrsim 300$. Moreover, the characteristic evolution time scale has been found to be well approximated by
$\tau=\frac{\Delta E}{4\pi\lambda^{2}V_{\pm}}.$ (17)
In all that follows we set $\Delta E=1$ and $V_{0}=3V_{-}$.
Finally, we write the initial state as
$|\psi(t_{0})\rangle=\sqrt{p_{-}(0)}|\psi_{-}\rangle+\sqrt{p_{0}(0)}|\psi_{0}\rangle+\sqrt{p_{+}(0)}|\psi_{+}\rangle.$ (18)
Here, $p_{x}(0)$ with $x\in\{-,0,+\}$ is the (a priori arbitrary) probability to find the system in macrostate $x$. Unless otherwise mentioned, the states $|\psi_{x}\rangle$ are normalized Haar randomly chosen states in the subspace ${\cal{H}}_{x}$ on which $\Pi_{x}$ projects. By choosing them Haar randomly we guarantee an unbiased choice, somewhat in spirit of a maximum entropy principle for pure states. In particular, given that $p_{x}(0)$ is the only available information, this allows us to ask about the typical behaviour of the system.
III.2 Results
Figure 3: Time evolution of $p_{-}(t)$ (dark blue), $p_{0}(t)$ (magenta) and $p_{+}(t)$ (cyan) over dimensionless time $t/\tau$ for an initial nonequilibrium state (a) or equilibrium state (b) for $D=10^{4}$ (solid lines) and $D=10^{2}$ (circles). The horizontal black lines indicate the equilibrium value $=(V_{-},V_{0},V_{+})/D$.
First, to get a feeling for the average dynamics, we consider the time evolution of the macrostate distribution $p_{x}(t)={\langle{\psi(t)|\Pi_{x}|\psi(t)}\rangle}$ in Fig. 3 for two different Hilbert space dimensions. In Fig. 3(a) we start from the nonequilibrium initial condition $[p_{-}(0),p_{0}(0),p_{+}(0)]=(1,0,0)$. For $D=10^{4}$ (solid lines) we see an exponential (Markovian) decay of the probabilities to their equilibrium value $(V_{-},V_{0},V_{+})/D=(0.2,0.6,0.2)$ (thin horizontal black lines) as predicted by statistical mechanics. The characteristic evolution time scale is set by $\tau$ and equilibration happens roughly for $t\gtrsim 7\tau$. For the smaller Hilbert space dimension of $D=10^{2}$ (circles) a similar tendency but much larger fluctuations are observed, as expected.
To continue with the many worlds simulation, we now choose an equilibrium initial state with $p_{x}(0)=V_{x}/D$, for reasons that will also become clear in Sec. IV. Naively, one would expect nothing interesting in such a Universe as indicated in Fig. 3(b): aside from fluctuations, which are exponentially suppressed as a function of the Hilbert space dimension, nothing seems to happen. But interestingly, a completely different picture is possible: even from an equilibrium state many different interesting nonequilibrium universes can emerge. However, to warrant such a conclusion, we first of all need to ensure that the histories defined by the present coarse-graining are decoherent, i.e., they obey the DHC condition as discussed in the previous section. In the following, the evaluation of the DF is done for constant time intervals $t_{k+1}-t_{k}$ equal to the nonequilibrium relaxation time scale $\tau$ defined in Eq. (17).
The emergence of decoherence is shown in Figs. 4 and 5. First, Fig. 4 plots the average off-diagonal elements of the normalized DF as defined in Eq. (5) for histories of lengths $L\in\{2,3,4,5\}$ as a function of the Hilbert space dimension $D$. Each marker in Fig. 4 corresponds to one particular random matrix Hamiltonian with one particular Haar random initial equilibrium state defined in Eq. (18). In total, we consider three different realizations of the random matrix Hamiltonian and, for each, three different realization of the initial state, amounting to $3\times 3=9$ different realizations. To extract the overall trend, we average the markers and fit a scaling law of the form $D^{-\alpha}$ (solid red line).
Figure 4: Plot of $\epsilon$, the average violation of decoherence defined in Eq. (5), in the weak coupling regime as a function of the Hilbert space dimension for histories of length $L=2$ (a), $L=3$ (b), $L=4$ (c) and $L=5$ (d). Blue crosses mark results for a single realization of the random matrix interaction and the random initial equilibrium state. Red solid lines fit a scaling law of the form $D^{-\alpha}$ to their averages and the value of $\alpha$ is indicated in the legend inset. Note the double logarithmic scale.
[…]
${\langle{\psi(\textbf{y})|\psi(\textbf{x})}\rangle}\sim\frac{1}{D}\sum_{i}c_{i}^{*}(\textbf{y})c_{i}(\textbf{x})\sim\frac{1}{\sqrt{D}}.$ (19)
While it seems plausible, it remains an open question whether this inverse square root dependence holds for all weakly perturbed random matrix theory models and Haar random initial equilibrium states. In any case, the scaling law suggests that quantum effects are exponentially suppressed as a function of the particle number $N$. Moreover, these conclusions are robust under changing the length $L\in\{2,3,4,5\}$ of the histories. Partial results for very long histories have been also recently obtained by two of the authors Strasberg and Schindler (2023).
Figure 5: Plot of $\Delta_{L}^{\text{max}}$, the worst case violation of decoherence defined in Eq. (11), with otherwise the same characteristics as in Fig. 4.
The same conclusions are also reached in Fig. 5, where we plot the maximum value of the trace distance defined in Eq. (11). Also for this more rigorous quantifier of decoherence we observe typicality and an exponential suppression of quantum effects as a function of the particle number $N$. However, since $\Delta_{L}^{\text{max}}$ is determined by the worst case scenario, stronger fluctuations are visible compared to $\epsilon$, where the average tames fluctuations. Moreover, the exponent $\alpha\approx 0.4$ is slightly smaller than the exponent extracted in Fig. 4. Nevertheless, Fig. 5 shows that also statistical outliers, which could be potentially detected by a clever and patient experimentalist, do not corrupt our conclusions. We conclude that the emergence of decoherence seems to be a stable and robust phenomenon.
Figure 6: (a) Depiction of the branching structure of the Multiverse with respect to the initial time $t_{0}$ and with respect to our chosen coarse-graining. (b) Histogram for the probabilities of different histories. For $L=3$ (c), $L=4$ (d) and $L=5$ (e), respectively, we plot the probability of universes with different (net) arrows of time as explained in the main text. All probabilities were obtained for a single realization of the interaction Hamiltonian (at weak coupling and for $D=10000$) and initial equilibrium state.
The picture that emerges from our model is depicted in Fig. 6. Figure 6 (a) shows a branching tree structure with respect to histories defined by energetic macrostates labeled by $x\in\{-,0,+\}$ for $L=3$ time steps. Importantly, we can now reason about these histories using a classical ontological model for sufficiently large $D$. To facilitate talking about it, we use Boltzmann’s entropy concept $S_{B}(x)=k_{B}\ln V_{x}$. We label transitions of the system from a lower to a higher Boltzmann entropy state (i.e., from $+$ or -- to 00) as “forward” transitions (blue dashed lines) to indicate that they comply with the conventional arrow of time when $t_{0}$ denotes the initial time. However, “backward” transitions (pink dash-dotted lines) from a higher to a lower entropy state (i.e., from 00 to $+$ or --) are also possible. Finally, transitions involving no change in Boltzmann entropy are labeled by “no arrow” (black lines). The Multiverse thus consists of different histories or branches that do not interfere and that describe universes with different arrows of time (including the possibility of universes with no arrow of time and both arrows of time) determined by the question whether heat flows from hot to cold or vice versa. For one realization of the Hamiltonian and one realization of the initial state we depict the probability for each history, which can be now interpreted classically, by the histogram below the tree-like branching structure in Fig. 6 (b). As a consistency check, we immediately see that the history $\textbf{x}=(0,0,0)$, in which the universe resides in the dominant equilibrium macrostate $x=0$, is the most probable one. Note that the probability for this history is different from the probability $(V_{0}/D)^{3}=0.6^{3}=0.216$ that one would expect by sampling the histories at time intervals comparable (or larger) to the equilibration time. The time steps $\tau$ we use are therefore clearly in the nonequilibrium regime.
Figure 7: Plot of $\epsilon$ for a random initial energy eigenstate. Blue crosses refer to three different randomly chosen eigenstates repeated for three different realizations of the random matrix coupling, i.e., nine realizations in total. Other details are as in Fig. 4.
An important consistency check is to ask about the probability of forward or backward arrows of time. Since the Schrödinger equation obeys time reversal symmetry and since the initial equilibrium state that we chose also does not introduce any time asymmetry, we should expect to get a time symmetric answer in this case—a point to which we return from a broader perspective in Sec. IV.4. The answer is shown in Fig. 6 (c), (d) and (e) where we plot the probability for a forward arrow of time (blue), no arrow of time (black) and a backward arrow of time (pink) for histories of length $L=3$, $L=4$ and $L=5$, respectively. Here, histories with multiple arrows of time contribute according to their “net” arrow of time, i.e., the history $\textbf{x}=(0,+,0)$ has no overall arrow of time. Owing to the fact that we have only three macrostates, there can be no histories with two or more net forward or backward arrows of time. To conclude, we see that the arrows are indeed symmetrically distributed in our model as it should be. Some non-visible small random fluctuations (in the third significant digit) are due to the fact that a random initial state can still have a small preference to evolve into a higher or lower entropy region, i.e., it is not perfectly symmetric under time reversal. The histograms in Fig. 6 are based on a single realization of the Hamiltonian and initial state, but we again found this behaviour to be typical.
We continue by challenging our model. This is first done by considering an atypical initial state in form of a randomly selected energy eigenstate $|k\rangle$ of the full Hamiltonian. From the eigenstate thermalization hypothesis Deutsch (1991); Srednicki (1994, 1999); D’Alessio et al. (2016); Deutsch (2018); Reimann and Dabelow (2021) it is known that these states give rise to probabilities $p_{x}={\langle{k|\Pi_{x}|k}\rangle}$ very close to the thermal (microcanonical) prediction $V_{x}/D$, i.e., a picture similar to Fig. 3 (b). However, it is currently not known whether multi-time correlation functions such as the DF display the same behaviour for energy eigenstates and Haar randomly sampled equilibrium states.
Figure 8: Plot of $\Delta_{L}^{\text{max}}$ with otherwise the same characteristics as in Fig. 7.
The results for the emergence of decoherence with initial energy eigenstates are shown in Figs. 7 and 8. We again observe a rather typical behaviour and a scaling law suggesting the exponential suppression of quantum effects as a function of the particle number. However, compared with Figs. 4 and 5 the variance in the datapoints is appreciably larger, i.e., there is less typicality. Moreover, the exponents are different. For $\epsilon$ we find $\alpha\approx 0.44$, which is close to the previous value of $\alpha\approx 0.5$. Instead, for $\Delta_{L}^{\text{max}}$ we find $\alpha\approx 0.2$, which is half of the value we found for Haar random initial equilibrium states. The preliminary conclusions that we can draw from these data is that correlation functions of energy eigenstates behave differently from Haar random equilibrium states (which will typically overlap with many energy eigenstates), which is an interesting result in its own right. In particular, Albrecht et al. recently introduced an “eigenstate einselection hypothesis” (Appendix B in Ref. Albrecht et al. (2022)), where they claim that the “general features of the consistent histories quantities are unchanged” for energy eigenstates. On a quantitative level we see here clear evidence for the opposite, but from a qualitative point of view we observe the same: the difference between $D^{-0.4}$ and $D^{-0.2}$ might not matter in practice, say, for $D\geq 10^{100}$. Thus, the emergence of decoherence as studied here seems to be a rather robust phenomenon if we restrict the attention to slow and coarse observables of a non-integrable quantum many-body system.
We remark that we have also studied the case of perturbed projectors of the form $V_{\delta}\Pi_{x}V_{\delta}^{\dagger}$, where the unitary $V_{\delta}=e^{iA\delta}$ rotates the projectors using a Hermitian matrix $A=A^{\dagger}$. However, for various $A$ we found quantitatively similar results (apart from fluctuations). [^5] We checked three cases: $A_{ij}=\delta_{i\pm 1,j}$ and $A_{ij}=\delta_{i+j,D+1}$ (with respect to the eigenbasis of the unperturbed Hamiltonian) and $A$ randomly drawn according to the same distribution as $H_{I}$. Therefore, and also to keep the manuscript concise, we do not display these results here, but we believe a systematic study is an interesting task for the future.
Instead, we find it important to illustrate that the emergence of decoherence is not a universal phenomenon valid for all observables. To demonstrate this, we consider the case of two strongly coupled systems exchanging energy. To this end, we increase $\lambda$ to $10\lambda$ such that the right hand side of the first equation in Eq. (16) equals one. Due to the strong interaction, local energy exchanges will now happen quickly. In some sense, it is no longer meaningful to talk about the local energies of $A$ and $B$ since the local energy levels of systems $A$ and $B$ will strongly hybridize to form new levels.
The result is a very different behaviour of decoherence as seen in Figs. 9 and 10. While typicality still holds well, the previously observed exponential suppression of quantum effects ceases to exist for histories of length $L\geq 3$. Instead, the scaling exponent of $\epsilon$ is around $\alpha=0.15$ in Fig. 9 and even negative values are possible for $\Delta_{L}^{\text{max}}$ as shown in Fig. 10 for $L=5$. Recalling that we have to base our conclusions on a finite (and rather small) amount of samples, a clear-cut conclusion seems not possible: exponential suppression of quantum effects, which requires $\alpha>0$, seems possible but is not warranted, and it is certainly much weaker than in the weak coupling (slow observable) regime.
Figure 9: Plot of $\epsilon$ at strong coupling with otherwise the same characteristics as in Fig. 4.
Notably, exponential suppression of quantum effects (with a large exponent $\alpha\approx 0.5$) still holds for the shortest length of histories with $L=2$. This seems counterintuitive, but it can be explained by recalling that we have chosen a Haar random initial equilibrium state of the form (18) with $p_{x}(0)=V_{x}/D$ in Figs. 9 and 10. The two-time DF then only probes quantum features of a two-time equilibrium correlation function. They are decoherent because, first, $\mathbb{E}[\Pi_{x_{0}}|\psi(t_{0}){\rangle\langle}\psi(t_{0})|\Pi_{y_{0}}]\sim\delta_{x_{0},y_{0}}$ where $\mathbb{E}[\dots]$ denotes a Haar random average, and second, typicality implies that most pure states $|\psi(t_{0})\rangle$ behave similarly. Indeed, for this situation an analytical proof of decoherence has been given in Ref. Strasberg et al. (2023).
We end this central part of the manuscript by pointing out that we have tested the robustness of our conclusions in various further situations for an initial Haar random equilibrium state. For instance, we considered random time spacings $t_{k+1}-t_{k}$ chosen uniformly either from $[0,\tau]$ or $[\tau,2\tau]$ (instead of the here considered constant time spacing of $t_{k+1}-t_{k}=\tau$), we considered random instead of equally spaced energies in the diagonal Hamiltonian blocks of Eq. (15), we used the Gaussian unitary ensemble instead of the Gaussian orthogonal ensemble for the off-diagonal blocks of Eq. (15), and we also plotted other quantifiers that one could derive from the DHC (3). In all these cases, we have observed the same qualitative behaviour, i.e., the typicality of a decay law of the form $D^{-\alpha}$. We do not show these results here to keep the manuscript at a reasonable length.
Figure 10: Plot of $\Delta_{L}^{\text{max}}$ at strong coupling with otherwise the same characteristics as in Fig. 5.
IV Consequences
For coarse and slow observables in isolated non-integrable many-body systems, the finite size scaling observed above shows an exponential suppression of quantum effects and that this suppression is typical, thus strongly suggesting that quantum interference effects vanish for macroscopic systems. Based on an arguably minimal set of assumptions this provides an explanation why the MWI is compatible with a classically perceived reality, provided we focus on the perception of slow and coarse observables (as we humans do). Experimentally, for any fixed measurement error the figures above can be used to interpolate at which point the coarse dynamics of an isolated quantum system becomes effectively indistinguishable from a classical stochastic process.
Cleary, there are many details to be filled in (see Sec. V), but in this section we take the perspective that the above theory is correct and explore its consequences. In particular, we consider from that perspective the set selection problem, the initial state of the Universe, the foundations of quantum mechanics and the arrow of time.
IV.1 Set selection problem
The set selection problem has been used to criticize the histories formalism by pointing out that the DHC is incomplete Paz and Zurek (1993); Dowker and Kent (1995, 1996); Riedel et al. (2016), see also Refs. Zurek (2003); Schlosshauer (2019); Zurek (2022). The basic reasoning behind the criticism is the following. Consider for fixed $L$ and $M$ the manifold H of all possible projectors that can be used to define histories x (of very different physical meaning). The dimension of this manifold scales like $D^{2L}$, whereas the number of constraints imposed by the DHC in Eq. (3) scales like $M^{2L}$. Thus, the submanifold DH of decoherent histories has a dimension that scales like $D^{2L}-M^{2L}$, which (even though of measure zero with respect to H) is enormous for $D\gg M$. Hence, which of the many physically very distinct histories should one use? [^6] Note the close similarity to the preferred basis problem.
One response to this problem is to claim that this is actually not a problem Griffiths (2019), but we would argue that there is no set selection problem if one focuses on slow and coarse observables of many-body systems (that is, situations relevant to human perception), and any remaining ambiguity is actually important. While this has been remarked in Ref. Griffiths (2019), a systematic explanation backed up by unambiguous evidence has not been given.
The crucial insight is to demand a certain stability or robustness of decoherence with respect to the initial state $|\psi(t_{0})\rangle$ and times $t_{k}$. To repeat: above we found approximate decoherence (without exception) for slow and coarse observables for a large set of initial states, times and Hamiltonians. In contrast, it seems likely that the vast majority of decoherent histories in DH is not robust in this sense. Instead, it likely depends sensitively on, e.g., $|\psi(t_{0})\rangle$, as sketched in Fig. 11.
For instance, one example is a coarse-graining $\{\Pi_{x}(t_{k})\}$ at time $t_{k}$ with one projector, say the first for $x=1$, equal to $\Pi_{1}(t_{k})=|\psi(t_{k}){\rangle\langle}\psi(t_{k})|$. Here, $|\psi(t_{k})\rangle=U_{k,0}|\psi(t_{0})\rangle$ is the unitarily evolved initial state. This coarse-graining satisfies by construction the DHC, but it is very sensitive to $\psi(t_{0})$ and $t_{k}$: changing them while leaving the projectors fixed will quickly destroy decoherence.
Figure 11: The square depicts the manifold H of all histories and the shade indicates the amount of decoherence (white regions correspond to no decoherence). The dark black line depicts the submanifold DH of exactly decoherent histores, which is very sensitive to different initial states $\psi(t_{0})$ (a) and $\psi^{\prime}(t_{0})$ (b). All $\texttt{DH}(\psi_{0})$ intersect for histories of conserved quantities (blue dot, exaggerated in size). Most important to our argument is the stability island (pink shaded region) of approximately decoherent histories, which remains (almost) unchanged for (a) and (b).
One might object that those examples are nevertheless legitimate, but there are at least two fundamental practical obstacles. First, it is unclear how humans should get access to a projector of the form $|\psi(t_{k}){\rangle\langle}\psi(t_{k})|$ using earthly instruments. Second, proper relativistic considerations teach us that at any given time $t$ we can only access part of the universal wave function $\rho_{\text{accessible}}(t)=\mbox{tr}_{\text{inaccesible}}\{|\psi(t){\rangle\langle}\psi(t)|\}$. Thus, even if we assume we find (by some magic instrument) a strange decoherent history depending on $\rho_{\text{accessible}}(t)$, decoherence would likely be destroyed in the next second once a new photon from an hitherto unobserved star reaches us.
Admittedly, this is not a rigorous argument showing that most of the histories in DH are of this fragile nature, but one should also admit in defense of this argument that an interesting example of stable and robust decoherent histories (beyond what we discussed here) has never been presented. Thus, we conjecture that the set selection problem is solvable by focusing on locally and practically accessible, approximately decoherent histories that are stable and robust.
We further strengthen the argument by pointing out the similarity with debates dating back to Boltzmann, Loschmidt and Zermelo. For a given initial state it is possible to find a continuum of observables that does not thermalize or entropies that violate the second law, a fact well known in pure state statistical mechanics. But all these artificially constructed counterexamples are of little relevance because the smallest perturbation of the initial state or Hamiltonian causes them to behave in agreement with thermodynamics again.
In addition, we believe that approximate (instead of exact) decoherence is key to resolve this problem because approximate decoherence can be very robust, see the “stability island” in Fig. 11.777In addition, our scaling law indicates for realistic many-body systems that approximate decoherence becomes practically indistinguishable from exact decoherence. Indeed, two different observers never perceive exactly the same observable (for instance, by reading of a display at slightly different angles), but both should still have compatible perceptions (this point is also central to quantum Darwinism Zurek (2009); Korbicz (2021); Zurek (2022)). Thus, instead of trying to find a single spot in the history space H that is exactly decoherent, it is more important to find the extended region that guarantees robust approximate decoherence. In this respect, we are sceptical about the idea that approximately decoherent histories can be always slightly distorted to become exactly decoherent Dowker and Kent (1995, 1996); Halliwell (2001, 2005): not only is it unclear how to define slight distortion, but it also seems unlikely that this distortion will be robust (independent of $\psi_{0}$ and $t_{k}$).
Thus, the picture that emerges is that of a robust stability island that describes accessible and almost exactly decoherent histories for large systems. This stability island is itself a huge region but for good reasons. First, it guarantees that different observers can have similar perceptions of classical reality. Second, the set of slow observables in this stability island will approximately commute for large systems, thus allowing one to define joint decoherent histories describing, e.g., the position, momentum and energy of a flying stone. [^8] However, the technical details of how to precisely handle approximately commuting observables are still subject to research Hastings (2009); Ogata (2013); Halpern et al. (2016).
We finish with two remarks. First, there is also an unappreciated set selection problem in EID. Gauge symmetries imply that there is a continuum of physically equivalent system-environment tensor product splittings of the same physical system Stokes and Nazir (2019, 2022). Each splitting gives rise to a different system-environment Hamiltonian and, consequently, a different pointer basis. Of course, what matters in practice is that the pointer basis is compatible with the experimental measurement procedure, but this resolution of the “pointer basis selection problem” is identical in spirit to the resolution proposed here in the histories context.
Second, we repeat that the proposed solution is restricted to slow and coarse observables of many-body systems that approximately decohere. It does not solve the set selection problem if one demands exact decoherence. Also other fundamental problems within the histories formalism such as the one debated in Refs. Kent (1997); Griffiths and Hartle (1998); Kent (1998) are not directly influenced by our argument.
IV.2 Initial state of the Universe
Our current understanding of the Universe is based on two separate concepts: dynamical laws (Newton, Maxwell, Schrödinger, Einstein, etc.) on the one hand, and boundary conditions—often in form of initial conditions—on the other hand. Therefore, much effort is devoted to understand the “state of the Universe”, often at some “initial” time where things were particularly simple.
This is true quantum mechanically and classically. For instance, the arrow of time and the second law suggest for a classical Universe that it must have started in an extremely exceptional and fine tuned low entropy state Feynman et al. (1963); Penrose (1989); Lebowitz (1993); Price (1996); Albert (2000), together with controversies as to whether this fine tuning requires further explanation or not Price (2004); Callender (2004). Moreover, the smoothness of the cosmic microwave background, as explained by the theory of inflation, suggests that quantum features have played a decisive role in the early Universe. Consequently, the wave function of the Universe plays an important role for various problems Hartle and Hawking (1983); Hawking (1984); Tipler (1986a); Mukhanov (2005); Carr (2007); Lehners (2023).
If the exponential suppression of quantum effects is typical, as predicted by the theory presented here, this suggests that the quest for the wave function of the Universe is more subtle than one might hope for, for two reasons. First, according to the MWI (together with an emergent classical structure) all we can infer is that the wave function of the Universe has some unknown but non-zero component compatible with our branch or history, an evident point made already before, e.g., in Ref. Tipler (1986b). Second, we also found that the emergence of decoherence is typical, i.e., almost all possible wave functions give rise to the same macroscopic Multiverse from a statistical point of view, i.e., up to random short fluctuations and unlikely statistical outliers. To see this, recall that our results were obtained for an initial equilibrium state and that one can mathematically rigorously show, using the concept of measure concentration (typicality), that almost all states are in equilibrium Farquhar and Landsberg (1957); Bocchieri and Loinger (1958); Gemmer et al. (2004); Popescu et al. (2006). Moreover, even when starting from an unlikely nonequilibrium macrostate, this state still lives in an enormously large subspace (according to Penrose’s estimate Penrose (1989) the entropy at the time of the big bang was around $10^{88}k_{B}$, leaving us still with on the order of ${\cal{O}}(10^{10^{88}})$ many microstates) and will typically give rise to decoherence.
Our results therefore imply that the emergence of decoherence requires no fine tuning. [^9] Note that we only refer to the fine tuning problem of the initial state. The fine tuning problem of the fundamental constants is not solved within the present considerations (see Ref. Carr (2007) for further discussion). Moreover, even a Universe in equilibrium gives rise to a Multiverse with many interesting nonequilibrium universes, even though their probability can be very small (we discuss this further in Sec. IV.4). Thus, at least without invoking further arguments, observing the fact that we can reason about our (local) universe classically implies almost nothing for the global wave function of the Universe.
IV.3 Foundations of quantum mechanics
We have just found that, as far as the emergence of decoherence is concerned, the MWI applied to slow and coarse observables gives rise to classical universes for almost all initial states. These initial states might be at equilibrium, very far from equilibrium (as the current big bang model suggests), or anything in between. Moreover, Sec. IV.1 argued that the histories framework, perhaps combined with other arguments such as quantum Darwinism, provides enough structure to reconcile the quantum Multiverse with our experience of living in one universe. It is thus scientifically questionable to reject the MWI only because of its seemingly “absurd” consequences. The MWI removes at least part of the fine tuning problem associated to the initial state of the Universe and it provides a basis for anthropic reasoning because there is something to select from (for a broad discussion see Ref. Carr (2007)).
These features of the MWI are not shared in any obvious way by interpretations of quantum mechanics that postulate a real objective (and not only apparent) collapse of the wave function due to measurements or hidden variables. Since these interpretations imply the existence of one single Universe, it seems likely that they face difficulties to explain the origin of our Universe without invoking a lot of lucky chance. This observation is interesting from a historical perspective because Ockham’s razor was a common objection against the MWI Shimony (1963): Why should one introduce a very large number of worlds that nobody can see? But not only can this argument be reversed—why and when should one introduce a large number of unnecessary collapses? Vaidman (2021)—but the MWI might even provide a simpler basis for cosmological considerations, see also Refs. Hartle and Hawking (1983); Hawking (1984); Tipler (1986a); Mukhanov (2005); Carr (2007); Lehners (2023).
IV.4 Arrow(s) of time
The Schrödinger equation obeys time reversal symmetry, and the initial equilibrium state used in the above simulation also does not introduce any time asymmetry. It is therefore an important consistency check Price (1996) that one finds (statistically) as many histories with a forward arrow of time as a backward one, as demonstrated in Fig. 6. Thus, arrows of time can be an emergent concept in a time reversal symmetric Universe (in varying contexts this point was also made in Refs. Carroll and Chen (2004); Barbour et al. (2014); Deutsch and Aguirre (2022)). Moreover, we remark that the DF can be as well defined with respect to the time-reversed evolution describing a branching of the wave function backwards in time (which would be natural for quantum retrodiction Barnett (2014)). Finally, there is actually also no preferred “initial” time $t_{0}$ in our model: we would observe the same phenomenology for any choice of $t_{0}$, a fact that is again related to our initial equilibrium state and the typicality of decoherence, see also Fig. 3 (b). Some special time $t_{0}$ could have been justified if one knows that the wave function of the Universe is at that time special too (i.e., out of equilibrium), but even then this would be a pure convention helpful for some calculations but not all.
In summary, our model demonstrates a time reversal symmetric system where arrows of time form on individual branches of the MWI. Unfortunately, owing to the restriction to only three macrostates $x\in\{-,0,+\}$, these arrows of time are minimal and do not suffice to explain our real world. While this problem can be overcome by considering more macrostates, branches further away from equilibrium will have less probability. This is a quantum version of the classical Boltzmann brain paradox (see also Ref. Albrecht and Sorbo (2004) for a discussion in context of inflation) that is often invoked to argue against the possibility that our universe is a mere fluctuation Carroll (2020), but a different reading is possible Müller (2020). Our model is still too simple to contribute to this debate, but it rigorously demonstrates at a quantitative level that classical universes with different symmetrically distributed arrows of time, even if they are mini-arrows, can emerge from equilibrium. This point is further discussed in Ref. Albrecht et al. (2022).
To better incorporate time reversal symmetry, one could modify the DF Isham (1994); Isham and Linden (1997); Griffiths (2002) or use other time symmetric formulations Aharonov et al. (1964); Vaidman (2010). Such a time symmetric view would then contradict common presentations of the MWI, where an exponentially increasing branching of the wave function occurs in one time direction (usually aligned with the authors’ own arrow of time). Indeed, some works went even further and identified this one-directional branching of the wave function with the increase in thermodynamic entropy Deutsch (2010); Aaronson (2013). In our view, there is nothing that warrants this conclusion, and it does not explain the apparent arrow of time.
V Concluding Perspectives
We presented a direct evaluation of the DF for multiple (up to five) time steps from first principles for a non-trivial example. We rigorously defined quantum interference effects and extracted a scaling law of the form $D^{-\alpha}$ by varying the Hilbert space dimension $D$ over four orders of magnitude. This provides a firm starting point to quantitatively discuss the DHC.
The resulting picture is that coarse and slow observables of non-integrable many-body systems give rise to a robust and stable form of approximate decoherence; see also Refs. Van Kampen (1954); Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Albrecht et al. (2022); Strasberg et al. (2023); Strasberg (2023) for supporting evidence. We have further argued that this robustness and stability could solve the preferred basis problem of the MWI or the set selection problem of the histories formalism, respectively.
Importantly, this picture does not contradict previous works, but it complements and extends them by providing new perspectives, tools and insights. For instance, our results do not rely on the widely used concepts of EID and quantum Darwinsim, but they are not in conflict with them. Indeed, a $1/\sqrt{D}$ scaling law has been also extracted for a random matrix model in the context of EID He and Wang (2014). Moreover, we believe non-integrability is a key factor, but it is usually not regarded as such in other works on the quantum-to-classical transition.
In fact, chaos has been even realized as detrimental for the emergence of classicality within the Wheeler-DeWitt equation due to the breakdown of the WKB approximaton and EID was invoked to remedy for that Calzetta and Gonzalez (1995); Cornish and Shellard (1998); Calzetta (2012). Within our non-relativistic quantum mechanical model we can not make any direct contribution to this question, but one view suggested by this work is that emergence of classicality is best viewed as a synergy of different mechanisms instead of a single all-ruling idea. Indeed, it has been recently pointed out that the emergence of classicality in the early Universe is still an unsolved puzzle Berjon et al. (2021), and the present perspective might be able to add a piece to it.
Besides providing a firm starting point to quantitatively discuss the DHC, our work also shows that it is possible to discuss a variety of further aspects quantitatively, for instance, the distribution of arrows of time, the number of Boltzmann brains or typicality aspects of the initial wave function of the Universe. More broadly speaking, our results improve the understanding of multi-time quantum processes, many-body physics and pure state statistical mechanics.
Also in these contexts many open questions remain. For instance, it would be desirable to get a better analytical understanding of the DHC. The results reported in Refs. Strasberg et al. (2023); Strasberg (2023) were restricted to three-time correlation functions and generalizing them to higher orders likely requires novel techniques (however, some progress for arbitrary $L$-time correlation functions in a different context was reported in Refs. Figueroa-Romero et al. (2019, 2021); Dowling et al. (2023a, b)). Further insights about the time scales at which decoherence arises are also desirable, and it is in particular also worth to ask about “recoherence times”: due to Poincaré recurrences the emergence of decoherence in a finite dimensional quantum system can not be a persistent, eternal phenomenon. In particular, it would be intriguing to find out how the recoherence time scales with the number of time steps $L$ or, more generally, with the net information aquired along a history. Whether non-integrability is really necessary for the robust emergence of decoherence from first principles (in particular without using ensemble averages) is another largely open question.
We further remark that we have used the conventional framework of non-relativistic quantum mechanics and assumed the validity of the Born rule. A significant fraction of research is devoted to interpreting or understanding the origin of the Born rule within the MWI (see, e.g., Refs. Vaidman (1998); Saunders et al. (2010); Aguirre and Tegmark (2011); Wallace (2012); Vaidman (2020); Zurek (2022) and references therein). The present approach could also add insights to this debate when considering very long histories for $L\gg 1$, as recently studied by two of the authors Strasberg and Schindler (2023).
Finally, our model calculations presently suggest that for any coarse and slow observable of a non-integrable system every possible classical history is realized. While this suggests the realization of a wide array of histories, it is not necessarily true that “everything that can happen will happen” (as sometimes portrayed in both scientific and popular accounts of the MWI). For instance, extremely unlikely (sequences of) outcomes might be restricted to very small subspaces because they are highly atypical, but our results indicate that large subspaces are needed for the emergence of classicality, see also Ref. Strasberg and Schindler (2023). Moreover, whether all the perceived randomness in our world stems from quantum effects Albrecht and Phillips (2014) or not Del Santo and Gisin (2019) remains a fascinating question too.
Acknowledgements
We are grateful to Anthony Aguirre, Josh Deutsch and Jonathan Halliwell for discussions. PS is financially supported by “la Caixa” Foundation (ID 100010434, fellowship code LCF/BQ/PR21/11840014) and acknowledges further support from the European Commission QuantERA grant ExTRaQT (Spanish MICINN project PCI2022-132965), by the Spanish MINECO (project PID2019-107609GB-I00) with the support of FEDER funds, the Generalitat de Catalunya (project 2017-SGR-1127) and by the Spanish MCIN with funding from European Union NextGenerationEU (PRTR-C17.I1). JS acknowledges support by MICIIN with funding from European Union NextGenerationEU (PRTR-C17.I1) and by the Generalitat de Catalunya.